# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

### v3

@@ 1 Introduction
Dataset Perception Reasoning Single view Multi $\dagger$ Video Task types QA quantity
@@
Table 1: Comparison of spatial understanding datasets. $\dagger$ Multi-view questions involve 3–5 images per QA pair.
We categorize spatial understanding into two levels: spatial perception, which encompasses fundamental spatial concepts such as depth, distance, relative positions, and camera viewpoints, representing low-level cognitive abilities; and spatial reasoning, which involves higher-order spatial understanding, requiring the ability to infer and construct 3D scenes implicitly. A critical aspect of this advanced ability is view matching, particularly in multi-view contexts.
Building on these two levels of spatial understanding, we introduce SPAR-7M, a dataset encompassing 33 tasks that cover a diverse range of complexities. These include basic perception tasks such as depth and distance estimation, more advanced reasoning tasks like cross-view object matching, and high-level spatial cognition tasks such as spatial imagination. To systematically organize these tasks, we classify them into three levels: low, medium, and high, reflecting a progression from Spatial Perception to Spatial Reasoning. To generate the necessary training data, we introduce a 2D data generation and annotation pipeline that derives rich supervision from 3D scene data. The resulting dataset, SPAR-7M, provides precise 3D ground-truth annotations, significantly improving both task diversity and reliability. Unlike existing approaches that rely solely on 2D annotations [10, 19], our dataset enables a more structured and accurate assessment of spatial understanding. Tab. 1 illustrates the diversity of input formats and task types within our dataset.
Despite the growing body of spatial benchmarks, many focus predominantly on high-level spatial reasoning, often overlooking fundamental spatial perception [24, 41, 40]. Furthermore, some benchmarks, particularly those using internet-collected images, struggle to effectively evaluate multi-view tasks, while others based on video inputs are not well-aligned with typical model inputs [52]. To address these limitations, we select 20 tasks from the SPAR-7M validation set. After carefully manual validation and filtering, we created SPAR-Bench, a comprehensive benchmark consisting of 7,207 questions that span tasks from basic perception to complex reasoning. SPAR-Bench supports both single-view and multi-view inputs, offering a more holistic and diverse evaluation of spatial understanding.
Our main contributions are as follows: (i) We introduce a 2D spatial QA generation pipeline leveraging 3D ground truth, enabling the creation of both single-view and multi-view data. Additionally, we present SPAR-7M, a comprehensive dataset covering a wide range of tasks from spatial perception to reasoning. (ii) We present SPAR-Bench, a benchmark for evaluating the spatial capabilities of existing models. (iii) By training models on both SPAR-7M and large-scale 2D internet datasets, we achieve state-of-the-art performance across multiple 2D spatial benchmarks. Moreover, even with 2D-only inputs, further fine-tuning on 3D task-specific datasets yields competitive results. We will open-source SPAR-7M and SPAR-Bench, facilitating further research in spatial understanding.
2 Related work
@@
Existing spatial understanding benchmarks primarily focus on high-level reasoning while often neglecting fundamental spatial perception tasks [41, 24, 40, 4, 42, 52]. Most of them rely on either single-view images or video inputs, with limited consideration of multi-view settings, despite their ability to better represent local 3D structures. Single-view images restrict the complexity of questions that can be formulated and provide an incomplete assessment of VLMs’ spatial reasoning capabilities. While some benchmarks use video inputs [52], many VLMs are not optimized for video processing, but this does not necessarily mean they lack spatial reasoning abilities. To address these limitations, we introduce SPAR-Bench, a more comprehensive benchmark that encompasses both fundamental and advanced spatial tasks. It supports both single-view and multi-view inputs, enabling a more systematic and reliable evaluation of spatial understanding.
@@ 3 Method
Our research aims to equip VLMs with robust and accurate spatial understanding, focusing on both fundamental 3D perception and higher-order spatial reasoning and imagination. To address the challenges posed by limited data, we propose a novel data generation pipeline that creates a diverse range of spatial understanding tasks. Our pipeline leverages 3D scenes from ScanNet [20], ScanNet++ [53], and Structured3D [55], ensuring a broad and diverse representation of indoor environments. Along with this, we introduce a new benchmark to provide a comprehensive evaluation of models’ spatial capabilities. In the following sections, we first detail our data generation pipeline (Sec. 3.1) and the key characteristics of SPAR-7M (Sec. 3.2). We then introduce SPAR-Bench (Sec. 3.3), followed by an analysis of data quality and the manual filtering process used to ensure benchmark reliability (Sec. 3.4). Additionally, in Sec. 3.5, we present a multi-view-based 3D grounding approach designed to enhance VLMs’ ability to tackle a broader range of 3D tasks.
@@
This section introduces our proposed data generation pipeline. The input consists of the scene mesh, 3D bounding boxes of objects, the intrinsic and extrinsic of camera sequences, and the image sequences. Through a series of steps, including filtering, extracting 3D information, and task construction, we generate diverse spatial understanding data. The key steps of this process are detailed below, with the overall workflow illustrated in Fig. 2.
@@
In the first step, we focus on reducing data redundancy by filtering out images with minimal camera movement. Specifically, each image is associated with a camera pose (position and orientation). If an image’s camera position is within a distance threshold of a retained image and its orientation differs by less than an angular threshold, it is considered redundant and removed. The whole process is similar to keyframing and the implementation details can be found in the appendix  6.2.
@@
In the second step, we extract global 3D information from the scene and create a unified data structure to support multi-view cross-retrieval and annotation. This structure facilitates dataset alignment and the generation of downstream tasks. Each scene is represented by two types of entities: Image records and object records.
  • Image record stores metadata for each image, such as its index, camera intrinsics and extrinsics, image dimensions, and the list of objects present. Each object is linked to its 3D bounding box, 3D center point in world coordinates, and its 2D bounding box on the image plane.
  • Object record contains global information for all objects in the scene, including their ID, category, and a list of images where they appear (referencing the corresponding image indices).
This structure creates a unified mapping between image and object data in world coordinates, simplifying the alignment and management of data from different sources. It also lays the foundation for multi-dataset annotations and task generation. Additional technical details can be found in the appendix 6.3.
@@
In the third step, we use the complete 3D scene information to automatically generate a variety of tasks and corresponding QA pairs. Each task is configured with key details such as task type (e.g., distance prediction, spatial relations), input image count, question format, and the object types involved (object-to-object or object-to-observer relations).
For each task, we follow a specific pipeline to generate annotations, ensuring flexibility to extend or customize tasks as needed. After generating the annotations, placeholders in the template (e.g., [label_A], [label_B], [left_right]) are replaced with actual object labels and computed spatial information. The templates are initially created by humans, expanded with the help of GPT, and then refined through manual selection to ensure diversity and accuracy.
While the process for generating tasks is consistent, each of the 31 tasks follows its own unique procedure based on its specific requirements. Detailed information about task pipelines, QA templates, and GPT prompting can be found in the appendix 6.4.
Figure 3: Dataset and benchmark visualization. We visualize some samples from our proposed SPAR-7M and SPAR-Bench. We show the diversity and richness of our dataset, which covers different task types, input types, and object types.
@@ 3.2 Spatial dataset construction
Based on our data generation pipeline (Sec. 3.1), we construct SPAR-7M, a large-scale dataset designed for spatial understanding. SPAR-7M consists of approximately 4K indoor 3D scenes from diverse sources, covering 33 task types with over 7 million QA pairs. Each QA pair is paired with camera intrinsics, extrinsics, and depth information, enabling both multi-view reasoning and 3D spatial inference. The dataset includes tasks ranging from basic spatial perception (e.g., depth estimation, distance measurement) to high-level spatial reasoning (e.g., camera position prediction, multi-view spatial relations) (Fig. 4).
To comprehensively evaluate different aspects of spatial understanding, we incorporate three types of input formats: single-view, multi-view, and video. Single-view tasks test a model’s ability to infer spatial properties from a single image, while multi-view tasks require integrating spatial information across different viewpoints. The video-based setting goes beyond individual frames to introduce higher-level spatial perception and scene-level reasoning, challenging the model to understand the spatial structure of a more complex environment. While not all tasks include every input format, this diversity ensures a broad assessment of spatial reasoning across various contexts.
In addition to varying input formats, we categorize tasks based on their spatial relationships into object-object (OO) and object-camera (OC) types. OO tasks focus on the relative positioning of objects, such as determining whether one object is to the left, right, or behind another. OC tasks, on the other hand, emphasize the relationship between objects and the camera, requiring models to infer viewpoint-dependent properties. For multi-image inputs, one image is randomly selected as the primary viewpoint, serving as the reference for spatial reasoning.
@@ 3.3 Spatial benchmark construction
To systematically evaluate spatial understanding, we introduce SPAR-Bench, a benchmark designed to assess models across diverse spatial reasoning tasks and input modalities. Unlike existing benchmarks that primarily focus on high-level reasoning or rely on video-based inputs, SPAR-Bench provides a more balanced evaluation by covering both spatial perception and reasoning in single-view and multi-view settings.
SPAR-Bench consists of 20 tasks covering depth estimation, distance prediction, spatial relation reasoning, and spatial imagination. These tasks are sampled from the SPAR-7M test set, with 400 QA pairs per task. After manual quality checking, the final benchmark includes 7207 QA pairs. Each task is formulated as either numerical prediction (fill-in-the-blank) or classification (multiple-choice selection), evaluated using Mean Relative Absolute Error (MRA) and accuracy, respectively.
To capture different levels of spatial understanding, single-view tasks focus on spatial inference from a single image, while multi-view tasks require reasoning across multiple perspectives, such as object matching and camera pose inference. A full list of tasks and details on data construction are provided in the appendix 6.1. We visualize the dataset and benchmark in  Fig. 3. More detailed QA of datasets can be found in appendix 6.5.
@@ 3.4 Data quality check
To ensure the reliability of our proposed benchmark, we conduct a human verification process. Specifically, we sample 400 questions per task in the validation set, forming an 8,000-question benchmark. These questions undergo manual inspection, taking approximately 140 human hours. After filtering out problematic cases, 7,207 questions were retained, resulting in a dataset acceptance rate of 90.1%.
@@
Figure 5: Our 3D grounding pipeline via VLM. For multi-view inputs, we decompose the task into frame selection and mono-3D grounding on the selected frame.
@@ 4 Experiments
Base VLM 32.4 74.2 78.5 46.6 58.3 76.5 2323 84.5 66.7 68.7
+EMOVA-2M 24.5 66.3 64.8 42.4 55.3 73.8 2186 80.2 63.1 63.8
+SPAR-mix 41.1 72.3 89.1 43.9 57.5 73.2 2163 79.9 64.7 62.9
Table 2: Comparison of model performance on various spatial understanding and 2D gerneral benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch 3 during training and inference.
@@
In this section, we assess the effectiveness of our dataset and investigate the potential upper limits of spatial reasoning in VLMs from three key perspectives:
  • Performance on existing spatial benchmarks: We examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks.
  • Evaluation on our in-domain benchmark SPAR-Bench: We assess the pretrained model’s capabilities using our proposed benchmark, SPAR-Bench, which provides a comprehensive evaluation of spatial understanding across various tasks.
  • Fine-tuning on 3D-specific tasks: To investigate the generalization ability of our approach, we perform supervised fine-tuning (SFT) on 3D-specific benchmarks, including spatial QA and 3D grounding tasks.
Furthermore, in Sec. 4.4, we analyze the impact of different data compositions and training strategies on model performance, offering insights into optimal configurations for enhancing spatial reasoning capabilities.
Pretraining.
To train a VLM with strong spatial capabilities while preserving generalization, we use a mixed pretraining dataset SPAR-mix. 40% of the data is uniformly sampled from our dataset, SPAR-7M, while the remaining 60% is uniformly sampled from general 2D data, sourced from the EMOVA configuration. This general data includes tasks such as image-text QA, chart understanding, and OCR. We pretrain the model on a total of 2 million samples, using the base model InternVL2.5-8B.
4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench [49].
The Cambrian Vision-Centric Benchmark (CV-Bench) is a large-scale dataset with 2,638 manually verified examples, designed to assess 2D spatial relationships and object counting. It repurposes standard vision datasets, formulating natural language questions to evaluate spatial reasoning. Visual prompts, such as bounding boxes, are used to reduce ambiguity and ensure reliable assessments.
VSI-Bench [52].
To evaluate the effectiveness of our dataset, we compare models pretrained on our SPAR-mix (a combination of our data and the EMOVA dataset) against those trained solely on 2M EMOVA dataset [12]. As shown in Tab. 2, Pretraining with our mixed dataset results in significant performance gains and achieves SOTA performance on both the CV-Bench 3D and VSI-Bench benchmarks, which are specifically designed to evaluate spatial reasoning capabilities.Moreover, compared to models trained exclusively on 2M EMOVA dataset, our approach maintains competitive 2D generalization performance, demonstrating that integrating our dataset effectively enhances 3D spatial reasoning while preserving 2D capabilities.
@@ 4.1 Evaluation on existing 2D spatial benchmarks
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet [20], ScanNet++ [53], and Structured3D [55], the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in Tab. 3. Notably, on NuScenes, an outdoor dataset, our approach yields an 18-point performance gain, demonstrating its strong generalization ability beyond indoor environments.

### v4

@@ 1 Introduction
We categorize spatial understanding into two levels: Spatial Perception, which encompasses fundamental spatial concepts such as depth, distance, relative positions, and camera viewpoints, representing low-level cognitive abilities; and Spatial Reasoning, which involves higher-order spatial understanding, requiring the ability to infer and construct 3D scenes implicitly.
Building on these two levels of spatial understanding, we introduce SPAR-7M, a dataset comprising 33 tasks that span a broad range of complexities. To facilitate systematic learning and evaluation, we organize all tasks into three levels—low, medium, and high—reflecting a natural progression from spatial perception to spatial reasoning. While previous datasets have predominantly focused on semantic descriptions, SPAR-7M places equal emphasis on numerical representations, enabling models to develop a more accurate and grounded understanding of 3D space. For example, we include fundamental perception tasks such as depth and distance estimation. We argue that under multi-view inputs, pixel- or feature-level matching is a foundational capability that supports higher-level spatial reasoning. Accordingly, many of our tasks—such as cross-view object matching and viewpoint transformation prediction—are explicitly designed to evaluate and strengthen this essential skill.
To generate the necessary training data, we introduce a 2D data generation and annotation pipeline that derives rich supervision from 3D scene data. The resulting dataset, SPAR-7M, provides precise 3D ground-truth annotations, significantly improving both task diversity and reliability. Unlike existing approaches that rely solely on 2D annotations [6, 7], our dataset enables a more structured and accurate assessment of spatial understanding.  Table 1 illustrates the diversity of input formats and task types within our dataset.
Dataset Perception Reasoning Single view Multi-view Video Task types QA quantity
@@
Table 1: Spatial understanding datasets comparison. Multi-view uses 3–5 images per QA pair.
Despite the growing body of spatial benchmarks, many focus predominantly on high-level spatial reasoning, often overlooking fundamental spatial perception [18, 19, 17]. Furthermore, some benchmarks, particularly those using internet-collected images, struggle to effectively evaluate multi-view tasks, while others based on video inputs are not well-aligned with typical model inputs [20]. To address these limitations, we select 20 tasks from the SPAR-7M validation set. After carefully manual validation and filtering, we created SPAR-Bench, a comprehensive benchmark consisting of 7,207 questions that span tasks from basic perception to complex reasoning. SPAR-Bench supports both single- and multi-view inputs, offering a more holistic and diverse evaluation of spatial understanding.
Our main contributions are as follows: (i). We present a 2D spatial QA generation pipeline using 3D ground truth to create both single- and multi-view data. This yields tasks covering numerical and semantic reasoning, forming a more comprehensive spatial understanding dataset. (ii). We present SPAR-Bench, a benchmark for evaluating the spatial capabilities of existing models. (iii). By training models on both SPAR-7M and large-scale 2D internet datasets, we achieve state-of-the-art performance across multiple 2D spatial benchmarks. Moreover, even with 2D-only inputs, further fine-tuning on 3D task-specific datasets yields competitive results.
2 Related works
@@
Existing spatial understanding benchmarks mainly focus on high-level reasoning while often neglecting fundamental spatial perception tasks [19, 18, 17, 13, 14, 20]. Most of them rely on either single-view images or video inputs, with limited consideration of multi-view settings, despite their ability to better represent local 3D structures. Single-view images restrict the complexity of the questions and thus provide an incomplete assessment of VLMs’ spatial reasoning capabilities. While some benchmarks use video inputs [20], many VLMs are not optimized for video processing, but this does not necessarily mean they lack spatial reasoning abilities. To address these limitations, we introduce SPAR-Bench, a more comprehensive benchmark that encompasses both fundamental and advanced spatial tasks. It supports both single- and multi-view inputs, enabling a more systematic and reliable evaluation of spatial understanding.
@@ 3 Method
Our research aims to equip VLMs with robust and accurate spatial understanding, focusing on both fundamental 3D perception and higher-order spatial reasoning and imagination. To address the challenges posed by limited data, we propose a novel data generation pipeline that creates a diverse range of spatial understanding tasks. Our pipeline leverages 3D scenes from ScanNet [27], ScanNet++ [28], and Structured3D [29], ensuring a broad and diverse representation of indoor environments. Along with this, we introduce a new benchmark to provide a comprehensive evaluation of models’ spatial capabilities. In the following sections, we first detail our data generation pipeline (section 3.1) and the key characteristics of SPAR-7M (section 3.2). We then introduce SPAR-Bench (section 3.3), followed by an analysis of data quality and the manual filtering process used to ensure benchmark reliability (section 3.4). Additionally, in section 3.5, we present a multi-view-based 3D grounding approach designed to enhance VLMs’ ability to tackle a broader range of 3D tasks.
@@
This section presents our data generation pipeline (detailed in fig. 2), which takes as input a scene mesh, 3D object bounding boxes, camera parameters, and image sequences. Through filtering, 3D information extraction, and task construction, we produce diverse spatial understanding data.
@@
In the first step, we focus on reducing data redundancy by filtering out images with minimal camera movement. Specifically, each image is associated with a camera pose. If an image’s camera position is within a distance threshold of a retained image and its orientation differs by less than an angular threshold, it is considered redundant and removed. The whole process is similar to keyframing and the implementation details can be found in the appendix 6.2.
@@
In the second step, we extract global 3D information from each scene and organize it into a unified data structure for cross-view retrieval and downstream annotation. This structure includes two types of entities: image records and object records, which together create a consistent mapping between image and object data in world coordinates. An image record stores image-level metadata, including index, camera parameters, resolution, and visible objects. Each object is linked to its 3D bounding box, 3D center in world coordinates, and 2D projection. An object record contains object-level metadata, including ID, category, and a list of associated image indices. This compact representation facilitates multi-dataset alignment and supports scalable task generation. Technical details are provided in appendix 6.3.
@@
In the third step, we use the complete 3D scene information to automatically generate tasks and corresponding QA pairs. Each task is configured with key details such as task type (e.g., distance prediction, spatial relations), input image count, question format, and the object types involved (object-to-object or object-to-observer relations). For each task, we follow a specific pipeline to generate annotations, ensuring flexibility to extend or customize tasks as needed. After generating the annotations, placeholders in the template (e.g., [label_A], [label_B], [left_right]) are replaced with actual object labels and computed spatial information. The templates are initially created by humans, expanded with the help of GPT, and then refined through manual selection to ensure diversity and accuracy. Detailed information about task pipelines, QA templates, and GPT prompting can be found in the appendix 6.4.
Figure 3: Dataset and benchmark visualization. We present illustrative samples from SPAR-7M and SPAR-Bench to highlight the diversity of spatial tasks and input types. The examples have been simplified for clarity; Full task formats are provided in the released dataset and Appendix 6.5.
@@ 3.2 Spatial dataset construction
Based on our data generation pipeline, we construct SPAR-7M, a large-scale dataset designed for spatial understanding. SPAR-7M consists of 4,000+ indoor 3D scenes from diverse sources, covering 33 task types with over 7 million QA pairs. Each QA pair is paired with camera parameters, and depth, enabling both multi-view reasoning and 3D spatial inference. The dataset includes tasks ranging from basic spatial perception to high-level spatial reasoning (detailed in fig. 4).
To evaluate diverse aspects of spatial understanding, we design tasks with three input formats: single-view, multi-view, and video. Single-view tasks assess spatial inference from one image, while multi-view tasks require cross-view integration. Video inputs introduce scene-level reasoning and long-range spatial perception. This diversity enables a comprehensive evaluation across settings. Tasks are further categorized by spatial relation type: object-object (OO) and object-camera (OC). OO tasks focus on inter-object positioning (e.g., left/right/behind), while OC tasks involve viewpoint-dependent relationships. For multi-view inputs, one image is randomly chosen as the primary viewpoint for reasoning.
@@ 3.3 Spatial benchmark construction
SPAR-Bench is constructed to evaluate models across a wide range of spatial tasks, from low-level geometric estimation to high-level relational reasoning. It includes 20 tasks sampled from the SPAR-7M test set, each with 400 QA pairs and manually verified for quality, resulting in a total of 7207 QA samples. The tasks span both single-view and multi-view settings and are formulated as either numerical prediction or classification, evaluated using Mean Relative Absolute Error (MRA) or accuracy, respectively. While the full dataset includes 33 spatial tasks, we select 20 for evaluation. Tasks requiring video input are excluded to disentangle spatial understanding from temporal modeling. Our focus is on evaluating how well models reason about 3D structure, geometry, and viewpoint relations from one or more views, independent of motion or continuity cues that video inputs may introduce. This also avoids overlap with existing video-based spatial benchmarks, and positions our evaluation as a complementary perspective focused purely on spatial reasoning.
To cover different facets of spatial understanding, the benchmark includes tasks such as depth estimation, distance comparison, object relation reasoning, and spatial imagination. Single-view tasks assess inference from a static observation, while multi-view tasks require reasoning across perspectives. A full task list is available in appendix 6.1, with QA visualizations in appendix 6.5, and dataset examples shown in fig. 3.
@@ 3.4 Data quality check
To ensure the reliability of our benchmark, we conduct a human verification process. Specifically, we sample 400 questions per task in the validation set, forming an 8,000-question benchmark. These questions undergo manual inspection, taking approximately 140 human hours. After filtering out problematic cases, 7,207 questions were retained, resulting in a dataset acceptance rate of 90.1%.
@@
The 90.1% acceptance rate indicates that the vast majority of samples were logically valid, clearly phrased, and visually answerable. In contrast, the 67.27 average human score table 4 reflects the intrinsic difficulty of the proposed spatial reasoning tasks—particularly those involving occlusion, viewpoint ambiguity, or subtle geometric cues. Rather than suggesting flaws in annotation quality, this gap underscores the challenging nature of our benchmark and its ability to expose limitations in both human and model spatial understanding.
Figure 5: Our 3D grounding pipeline via VLM.
@@ 4 Experiments
In this section, we assess the effectiveness of our dataset and investigate the potential upper limits of spatial reasoning in VLMs from the following perspectives: (i). Performance on existing spatial benchmarks: we examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks; (ii). Evaluation on our in-domain benchmark SPAR-Bench: we examine how pretraining on SPAR-7M improves model performance on image-based spatial benchmarks; (iii). Fine-tuning on 3D-specific tasks: To investigate the generalization of our approach, we perform supervised fine-tuning (SFT) on 3D-specific benchmarks, including spatial QA and 3D grounding tasks.
Furthermore, in section 4.4, we analyze the impact of different data compositions and training strategies on model performance, offering insights into optimal configurations for enhancing spatial reasoning capabilities.
Pretraining.
To train a VLM with strong spatial capabilities while preserving generalization, we use a mixed pretraining dataset SPAR-mix. 40% of the data is uniformly sampled from our dataset, SPAR-7M, while the remaining 60% is uniformly sampled from general 2D data, sourced from the EMOVA configuration. This general data includes tasks such as image-text QA, chart understanding, and OCR. We pretrain the model on a total of 2 million samples, using the base model InternVL2.5-8B. Pretraining was performed on 64 NPUs for 4 days, resulting in a total of about 6,100 NPU-hours.
4.1 Evaluation on existing 2D spatial benchmarks
CV-Bench [33].
The Cambrian Vision-Centric Benchmark (CV-Bench) is a 2,638-sample dataset for evaluating 2D spatial reasoning and object counting. It repurposes vision datasets with natural language questions and uses visual prompts (e.g., bounding boxes) to reduce ambiguity and ensure reliable evaluation.
VSI-Bench [20].
To evaluate the effectiveness of our dataset, we compare models trained on SPAR-mix (ours with EMOVA-2M) against those trained solely on EMOVA-2M. Since our dataset focuses specifically on spatial perception and reasoning, we incorporate a subset of QA data from EMOVA to maintain performance on general-purpose vision-language tasks. Note that the Base VLM has already been pretrained on extensive data with optimized training strategies, making it an unfair baseline for assessing the incremental gain. Therefore, we consider Base VLM + EMOVA-2M as the proper baseline. These results validate the effectiveness of our dataset design and indicate that spatial supervision can be introduced without compromising general QA performance. As shown in table 2, our model achieves state-of-the-art results on CV-Bench 3D and VSI-Bench, demonstrating significant improvements in spatial understanding. While there are minor drops on general benchmarks such as MMBench and TextVQA—mostly within 1%—this is expected, as our data emphasizes spatial reasoning rather than fine-grained vision-language alignment.
@@ 4.1 Evaluation on existing 2D spatial benchmarks
Base VLM 32.4 74.20 78.50 46.61 58.33 76.53 2323 84.45 66.67 68.73
+EMOVA-2M 24.5 66.27 64.83 42.40 55.25 73.8 2186 80.24 63.14 63.78
+SPAR-mix 41.1 72.25 89.08 43.92 57.48 73.2 2163 79.90 64.71 62.91
Table 2: Comparison of model performance on various spatial understanding and 2D general benchmarks. The base VLM is InternVL2.5-8B, and we use dynamic patch of 3. Note: Our main comparison is between +EMOVA-2M and +SPAR-mix, as both build upon the same Base VLM and general QA data. For detailed discussion, see section 4.1.
@@
We also find that pretraining with our dataset enhances the model’s spatial understanding across diverse, non-homogeneous datasets. Although our dataset is primarily derived from indoor scene datasets such as ScanNet [27], ScanNet++ [28], and Structured3D [29], the pretrained model shows consistent improvements across all tasks constructed with datasets different from ours in CV-Bench 3D, as shown in table 3.
Notably, our method achieves an 18-point gain on the outdoor nuScenes dataset, highlighting strong generalization beyond indoor scenes. This improvement stems from the spatial reasoning skills emphasized in our dataset—many CV-Bench 3D questions focus on relative distance comparisons, similar to tasks like DistI-OO and ObjRel. Since these tasks rely on 3D geometric cues rather than appearance or domain-specific semantics, the learned spatial priors transfer effectively across diverse environments.